Can anyone help me? I swear it's the last one:-)My problem is: Use the equations of the lines to determine if they're parallel. y=0.15x+0.79, 0.11x-y=o.85. I would be greatful if someone could show me how this is done, no matter how much I read and study, I still don't get it. Again Than-x and Have a Good Easter.
Answer by Alan3354(69443)

Use the equations of the lines to determine if they're parallel. y=0.15x+0.79, 0.11x-y=o.85.
----------------
Find the slopes of the 2 lines by putting the eqns in the slope-intercept form, y = mx+b. m is the slope.
-----------------
All you have to do it solve for y.
y=0.15x+0.79 eqn 1
0.11x-y=o.85 --> y = 0.11x - 0.85 eqn 2
m1 = 0.15
m2 = 0.11
The slopes are different, so they're not parallel.